Test the null hypothesis H0:μ=3.2against the alternative hypothesis HA:μ<3.2, based on a random sample of 25 observations drawn from a normally distributed population with x¯=3 and σ=0.71.

a) What is the value of the test statistic?

Round your response to at least 3 decimal places.

   

b) What is the appropriate p-value?

Round your response to at least 3 decimal places.

   

c) Is the null hypothesis rejected at:

i) the 10% level of significance?      NoYesClick for List  

ii) the 5% level of significance?      NoYesClick for List  
Answer #1

Solution hypothesis is The Noll and to M = 3,2 Haigu <3.2 a n test =0.71 - 25 statistic = a de-M - 3-3.2 0.71 25 z =-1.408 b)